OpenStudy (anonymous):

A cone has a base area of 12 in² and a lateral area which is 3.5 times bigger than the base area. The surface area of the cone is is it 42 ? if not what is it and whats the formula

OpenStudy (lgbasallote):

SURFACE AREA = Lateral Surface Area + Base Area

OpenStudy (lgbasallote):

so i don;t think your answer is right

OpenStudy (kropot72):

Base area = 12 sq.in. lateral area = 12 * 3.5 = 42 sq.in Total area = base area + lateral area = 12 + 42 = 54 sq.in.

OpenStudy (anonymous):

im always off by a bit thanks kropot72

OpenStudy (kropot72):

You're welcome :) Don't worry, you got part of the way there!